FOR THE same price you can order a basic pizza and add any two toppings from a choice of six different toppings. If a double of one topping is allowed, the number of different pizzas with two toppings is.
a. 36
b. 30
c. 12
d. 6

Answer by stanbon(75887)

FOR THE same price you can order a basic pizza and add any two toppings from a choice of six different toppings. If a double of one topping is allowed, the number of different pizzas with two toppings is.
-------------------------------------
Pick a basic with two toppings: 6C2 = 6*5/1*2 = 15
Pick a double of a topping or do not: 2 choices
Ans: 2*15 = 30
-------------------------
Cheers,
Stan H.
-------------------------
a. 36
b. 30
c. 12
d. 6